# Bayesian Concept Bottleneck Models with LLM Priors

### v1

@@ Appendix E Example prompt for annotations for candidate concepts (Step 3)
Appendix F Experiment details
Note on computation time: When running BC-LLM with GPT-4o-mini-2024-07-18 and GPT-4o, each iteration completed within a few minutes.
F.1 MIMIC
@@
[…]
One-pass summarization and BC-LLM were run to fit CBMs with $K=6$ concepts. BC-LLM was run for $T=5$ epochs. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration.
F.2 CUB-Birds dataset
To train a bird classifier for $R$ subtypes, One-pass summarization and BC-LLM were run to fit CBMs with $K=\min(10,\max(4,R))$ concepts. BC-LLM was run for a maximum of $T=5$ epochs or 25 iterations, whichever was reached first. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration.
F.3 Clinical notes and tabular data from ZSFG
To predict readmission risk, the models were trained to analyze the sections “Brief history leading to hospitalization” and the “Summary of hospitalization” in the discharge summary for each patient. One-pass summarization and BC-LLM were run to fit CBMs with $K=4$ concepts. BC-LLM was run for $T=5$ epochs. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration. For the black-box comparator, we used a ResNet50 model pre-trained on ImageNetV2, which ensures ResNet’s training set does not overlap with the CUB-Birds dataset.
For the OOD experiment, the Bunting CBM was applied to 30 images of finches.
Appendix G Laplace Approximation of Split-Sample Posterior
@@
Remark G.1.
The MAP estimate $\vec{\theta}_{\operatorname{MAP}}$ can be computed by a call to scikit-learn’s LogisticRegression() model, setting C $=2\gamma^{2}$.
Remark G.2.
@@
Appendix H Proof of Theorem 3.1
@@
This quantifies how relevant the concepts in $\vec{c}$ are to the prediction of the response $Y$. Note that the definition does not require that the model class is correctly specified (i.e. there does not have to be some $\vec{c}$ and $\vec{\theta}$ so that (1) holds). Before presenting the proof, we first detail the assumptions we make.
Assumption H.1.
  1. (i)
@@
  6. (vi)
    Let $\mathcal{C}_{n}\coloneqq\cup_{\vec{c}_{-k},\mathbf{y},\mathbf{X}}\{c\colon Q(c;\vec{c}_{-k},\mathbf{y},\mathbf{X})>0\}$. Assume $\mathcal{C}_{n}$ is a finite set of size at most $\exp(n^{1-\epsilon})$ for some $10<\epsilon<$ 1;
  7. (vii)
@@
Step 2: Reduction to connected components. Since $(\tilde{\vec{c}}^{~{}t})_{t=1}^{\infty}$ satisfies the detailed balance equations, it is a symmetric Markov chain, i.e. it has an equivalent representation as a flow on a network, in which the vertices are states and the edges correspond to nonzero transition probabilities. This property is inherited by $(\vec{c}^{~{}t})_{t=1}^{\infty}$. Consider a connected component $\mathcal{A}$ of the network. The Markov chain, when restricted to $\mathcal{A}$, is aperiodic, since there must exist a self-loop. Furthermore, it is finite according to Assumption H.1(vi). As such, $(\vec{c}^{~{}t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ (LevinPeresWilmer2006). Indeed, all stationary distributions of $(\vec{c}^{~{}t})_{t=1}^{\infty}$ are convex combinations of such distributions. Hence, it suffices to show the desired property for $\pi_{\mathcal{A}}$.
Step 3: Concentration. In order to analyze $\pi_{\mathcal{A}}$, we need quantitative control over the transition probabilities via concentration. For a given dataset $\mathcal{D}$ of size $n$, we condition on the event guaranteed by Proposition H.3.
@@
so long as the denominator on the right hand side is nonzero. To see this, notice that any state $j$ for $b+1\leq j\leq a$ corresponds to a vector $\tilde{\vec{c}}^{~{}(0)}\notin\mathcal{C}^{*}$. However, using Assumption H.1(vii), there is a sequence $(\tilde{\vec{c}}^{~{}(0)},\tilde{\vec{c}}^{~{}(1)},\ldots,\tilde{\vec{c}}^{~{}(K-1)})$ such that for $k=1,\ldots,K$, we have (i) $\tilde{\vec{c}}^{~{}(k)}_{-k}=\tilde{\vec{c}}^{~{}(k-1)}_{-k}$, (ii) $Q(\tilde{c}_{k};\tilde{\vec{c}}^{~{}(k-1)}_{-k},\mathbf{y}_{S},\mathbf{X})\geq\eta$, (iii) $L(\tilde{\vec{c}}^{~{}(k)})\geq L(\tilde{\vec{c}}^{~{}(k-1)})$ with equality holding iff $\tilde{\vec{c}}^{~{}(k)}=\tilde{\vec{c}}^{~{}(k-1)}$. Combining (iii) with (32), we see that all acceptance ratios along this sequence are equal to 1, which means that the transition probability from state $j$ to states in $\mathcal{C}^{*}$ is at least
@@
The sum of entries in row $j-b$ in $P_{22}$ is the total probability of transition from state $j$ to other states within $\mathcal{A}\backslash\mathcal{C}^{*}$ and hence has value less than $1-\eta^{K}$. This holds for any row. Applying Lemma H.2, we thus get $\|P_{22}\|_{op}\leq 1-\eta^{K}$, which allows us to bound the denominator in (24). Next, every entry in $P_{12}$ is the probability of transition from a state in $\mathcal{C}^{*}$ to a state in $\mathcal{A}\backslash\mathcal{C}^{*}$. Using Step 3 and Assumption H.1(v), the acceptance probability $\alpha$ of such a transition satisfies
@@
Proposition H.3 (Concentration of split-sample posterior).
Assuming the conditions in Assumption H.1, then for any $\epsilon^{\prime}<\epsilon/2$, we have
@@
We can thus use Talagrand’s comparison inequality (vershynin2018high) to get, for any $t>0$, that
@@
[…]
@@
[…]
@@
Appendix I Stationary Distribution Under Consistent Proposals
@@
Proposition I.1.
Assuming equality holds in (2), the Markov chain defined by running Gibbs sampling (Algorithm 1) with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update has the posterior $p(\vec{C}|\mathbf{y},\mathbf{X})$ as a stationary distribution.
@@
as we wanted.
To show the statement for Multi-SS-MH-Update, we use its equivalence, for a fixed $S$, to the modified Multiple-Try Metropolis-Hastings method described in Appendix A. In Appendix A, we prove that this method satisfies the detailed balance equations for the posterior. Since the detailed balance equations are linear in the transition probabilities, we can marginalize this over $S$ to show that the posterior is stationary with respect for Multi-SS-MH-Update, which draws $S$ uniformly at random. ∎

### v2

@@ C.3 Example prompt for annotations for candidate concepts (Step 3)
Note that in cases where the LLM was unsure about the concept’s value, it was allowed to return a probability, which prior works have found to be helpful [6]. In general, the LLM outputted 1’s and 0’s. Probabilities were outputted occasionally, which as highlighted in the main manuscript, can be audited and checked by a human.
Appendix D Additional Results
D.1 MIMIC
We evaluate BC-LLM against baseline methods using Cohere’s Command-R LLM using the same settings as discussed in Section 4.2. Performance can be found in Figure 5.
Figure 5: Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts. Error bars indicate the 95% CI.
D.2 CUB-Birds dataset
We evaluate BC-LLM against baseline methods using Anthropic’s Claude 3.5 Haiku using the same settings as discussed in Section 4.1. Performance can be found in Table 2.
Method Accuracy ($\uparrow$) AUC ($\uparrow$) Brier ($\downarrow$)
BC-LLM 0.665 (0.593, 0.736) 0.849 (0.811, 0.888) 0.443 (0.367, 0.519)
LLM+CBM 0.583 (0.517, 0.648) 0.771 (0.722, 0.821) 0.507 (0.434, 0.580)
Boosting LLM+CBM 0.640 (0.568, 0.712) 0.811 (0.769, 0.852) 0.552 (0.455, 0.649)
Table 2: Results for the CUB-Birds dataset using Claude 3.5 Haiku
Appendix E Experiment details
E.1 Evaluating correct concept matches
Quantifying a “correct” concept match requires going beyond exact string matching, as there are many possible refinements for a single concept. For instance, for the concept “Does the note mention the patient smoke at present?”, there are rephrasings (“Does this note mention if the patient currently smokes?”), polar opposites (“Does the note not mention the currently smoking?”), or closely overlapping concepts (“Does the note mention the patient’s smoking history?”). To quantify if a learned concept was correct, we used the following rule: if the absolute value of the Pearson correlation for LLM annotations of two concepts exceeds 50%, the two concepts are sufficiently similar. We manually confirmed that any pair of learned and true concepts that passed this rule were always closely related.
E.2 MIMIC
@@
$\log{\frac{\Pr(Y=1|X)}{\Pr(Y=0|X)}}=$ $4*\mathbbm{1}\{\text{Does the note imply the patient is unemployed?}\}$
$+4*\mathbbm{1}\{\text{Does the note imply the patient is retired?}\}$
@@
LLM+CBM summarization and BC-LLM were run to fit CBMs with $K=6$ concepts. BC-LLM was run for $T=5$ epochs. Boosting LLM+CBM ran for 10 iterations, in which at most 1 new candidate was added each iteration.
E.3 CUB-Birds dataset
To train a bird classifier for $R$ subtypes, LLM+CBM summarization and BC-LLM were run to fit CBMs with $K=\min(10,\max(4,R))$ concepts. BC-LLM was run for a maximum of $T=5$ epochs or 25 iterations, whichever was reached first. Boosting LLM+CBM ran for 10 iterations, in which at most 1 new candidate was added each iteration. For the black-box comparator, we used a ResNet50 model pre-trained on ImageNetV2, which ensures ResNet’s training set does not overlap with the CUB-Birds dataset.
E.4 Clinical notes and tabular data from Hospital Anonymous
Data in this experiment contained PHI, for which IRB approval was obtained. To predict readmission risk, the models were trained to analyze the sections “Brief history leading to hospitalization” and the “Summary of hospitalization” in the discharge summary for each patient. LLM+CBM summarization and BC-LLM were run to fit CBMs with $K=4$ concepts. BC-LLM was run for $T=5$ epochs. Boosting LLM+CBM ran for 10 iterations, in which at most 1 new candidate was added each iteration.
E.5 Clinician survey
To conduct the survey for clinical relevance of features and concepts, the following instructions were given to clinicians:
Please rate the features learned by BC-LLM and its comparator methods. In the attached CSV, we have listed 25 or so features from various methods, where we mask which method has learned which feature. Please fill in the column "How clinically relevant is the candidate concept with determining a patient’s readmission risk? Enter values = 1: low, 2:medium, 3: high." Please try to use the full range of scores, rather than assigning all the features the same score.
Table 3 shows the table that clinicians were asked to fill in. The ordering of features was randomly shuffled to obfuscate which method had learned which concepts.
Table 3: Clinicians were asked to fill in the following table with their ratings of the clinical relevance of each feature/concept for predicting a patient’s readmission risk.
Candidate concept Score (1-3)
Does the patient have multiple medications prescribed?
Does the patient have a history of hypertension (HTN)?
Is this patient experiencing chest pain?
What is the patient’s weight?
Does the patient have heart failure with reduced ejection fraction?
What are the patient’s lab results for BNP?
Does the patient have substance dependence?
Is the patient stable at discharge?
Does the patient have a history of falls?
How many emergency department encounters does this patient have?
What are the patient’s lab results for Lactate Dehydrogenase?
Does the patient have uncontrolled diabetes?
Does the patient have a history of frequent hospital admissions?
How many encounters does this patient have?
Does the patient have a history of diabetes mellitus type 2 (DM2)?
Does the patient have complex medical history?
Does the patient have a history of sepsis?
What are the patient’s lab results for Creatinine?
Does the patient have a history of drug or substance use disorder?
What is the patient’s value for Expiratory Positive Airway Pressure (EPAP)?
Does the patient have had any missed outpatient appointments?
What are the patient’s lab results for Glucose?
Are the patient’s outpatient follow-up appointments scheduled?
Appendix F Laplace Approximation of Split-Sample Posterior
@@
Remark F.1.
The MAP estimate $\vec{\theta}_{\operatorname{MAP}}$ can be computed by a call to scikit-learn’s LogisticRegression() model, setting C $2\gamma^{2}$.
Remark F.2.
@@
Appendix G Proof of Theorem 3.1
@@
This quantifies how relevant the concepts in $\vec{c}$ are to the prediction of the response $Y$. Note that the definition does not require that the model class is correctly specified (i.e. there does not have to be some $\vec{c}$ and $\vec{\theta}$ so that $p\left(Y=1|X=x,\vec{\theta},\vec{c}\right)=\sigma\left(\sum_{k=1}^{K}\theta_{k}\phi_{c_{k}}(x)+\theta_{0}\right)$). Before presenting the proof, we first detail the assumptions we make.
Assumption G.1.
  1. 1.
@@
  6. 6.
    Let $\mathcal{C}_{n}\coloneqq\cup_{\vec{c}_{-k},\mathbf{y},\mathbf{X}}\{c\colon Q(c;\vec{c}_{-k},\mathbf{y},\mathbf{X})>0\}$. Assume $\mathcal{C}_{n}$ is a finite set of size at most $\exp(n^{1-\epsilon})$ for some $0<\epsilon<1$;
  7. 7.
@@
Step 2: Reduction to connected components. Since $(\tilde{\vec{c}}^{~t})_{t=1}^{\infty}$ satisfies the detailed balance equations, it is a symmetric Markov chain, i.e. it has an equivalent representation as a flow on a network, in which the vertices are states and the edges correspond to nonzero transition probabilities. This property is inherited by $(\vec{c}^{~t})_{t=1}^{\infty}$. Consider a connected component $\mathcal{A}$ of the network. The Markov chain, when restricted to $\mathcal{A}$, is aperiodic, since there must exist a self-loop. Furthermore, it is finite according to Assumption G.1(vi). As such, $(\vec{c}^{~t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ [55]. Indeed, all stationary distributions of $(\vec{c}^{~t})_{t=1}^{\infty}$ are convex combinations of such distributions. Hence, it suffices to show the desired property for $\pi_{\mathcal{A}}$.
Step 3: Concentration. In order to analyze $\pi_{\mathcal{A}}$, we need quantitative control over the transition probabilities via concentration. For a given dataset $\mathcal{D}$ of size $n$, we condition on the event guaranteed by Proposition G.3.
@@
so long as the denominator on the right hand side is nonzero. To see this, notice that any state $j$ for $b+1\leq j\leq a$ corresponds to a vector $\tilde{\vec{c}}^{~(0)}\notin\mathcal{C}^{*}$. However, using Assumption G.1(vii), there is a sequence $(\tilde{\vec{c}}^{~(0)},\tilde{\vec{c}}^{~(1)},\ldots,\tilde{\vec{c}}^{~(K-1)})$ such that for $k=1,\ldots,K$, we have (i) $\tilde{\vec{c}}^{~(k)}_{-k}=\tilde{\vec{c}}^{~(k-1)}_{-k}$, (ii) $Q(\tilde{c}_{k};\tilde{\vec{c}}^{~(k-1)}_{-k},\mathbf{y}_{S},\mathbf{X})\geq\eta$, (iii) $L(\tilde{\vec{c}}^{~(k)})\geq L(\tilde{\vec{c}}^{~(k-1)})$ with equality holding iff $\tilde{\vec{c}}^{~(k)}=\tilde{\vec{c}}^{~(k-1)}$. Combining (iii) with (30), we see that all acceptance ratios along this sequence are equal to 1, which means that the transition probability from state $j$ to states in $\mathcal{C}^{*}$ is at least
@@
The sum of entries in row jbj-b in $P_{22}$ is the total probability of transition from state $j$ to other states within $\mathcal{A}\backslash\mathcal{C}^{*}$ and hence has value less than $1-\eta^{K}$. This holds for any row. Applying Lemma G.2, we thus get $\|P_{22}\|_{op}\leq 1-\eta^{K}$, which allows us to bound the denominator in (22). Next, every entry in $P_{12}$ is the probability of transition from a state in $\mathcal{C}^{*}$ to a state in $\mathcal{A}\backslash\mathcal{C}^{*}$. Using Step 3 and Assumption G.1(v), the acceptance probability $\alpha$ of such a transition satisfies
@@
Proposition G.3 (Concentration of split-sample posterior).
Assuming the conditions in Assumption G.1, then for any $\epsilon^{\prime}<\epsilon/2$, we have
@@
We can thus use Talagrand’s comparison inequality [56] to get, for any $t>0$, that
@@
$\begin{split}p(\mathbf{y}|\vec{c},\mathbf{X})&=\int p(\mathbf{y}|\vec{\theta},\vec{c},\mathbf{X})\frac{1}{(2\pi\gamma^{2})^{K/2}}\exp(-\|\vec{\theta}\|^{2}/2\gamma^{2})d\vec{\theta}\\ &\leq\exp(nL(\vec{c})+O(n^{1-\epsilon^{\prime}})).\end{split}$ (39)
@@
$\int p(\mathbf{y}|\vec{\theta},\vec{c},\mathbf{X})\frac{1}{(2\pi\gamma^{2})^{K/2}}\exp(-\|\vec{\theta}\|^{2}/2\gamma^{2})d\vec{\theta}\geq\frac{r^{K}\operatorname{Vol}(B(1))}{(2\pi\gamma^{2})^{K/2}}\exp(-2R^{2}/\gamma^{2})\exp\left(n(L(\vec{\theta}^{*})-\sqrt{K}r^{2}-O(n^{-\epsilon^{\prime}}))\right).$ (43)
@@
Appendix H Stationary Distribution Under Consistent Proposals
@@
Proposition H.1.
Suppose the LLM proposes from the conditional partial posterior distribution, i.e. $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$. Then the Markov chain defined by running Gibbs sampling (Algorithm 1) with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update has the posterior $p(\vec{C}|\mathbf{y},\mathbf{X})$ as a stationary distribution.
@@
as we wanted. ∎